OpenStudy (anonymous):

Show that f(n) = (n^3+2n)/ 2n is in O(n2).

OpenStudy (zarkon):

I assume you want this for \[n\to\infty\]

OpenStudy (zarkon):

and that it is \[O(n^2)\]

OpenStudy (anonymous):

yes sorry I missed the ^

OpenStudy (zarkon):

show that \[\lim_{n\to\infty}\left|\frac{(n^3+2n)/(2n)}{n^2}\right|<\infty\]

OpenStudy (anonymous):

why is it over n^2? b/c of O(n^2)? do you know what that stands for?

OpenStudy (anonymous):

I simplified it to (1/2) + (1/n^2)

OpenStudy (anonymous):

so at infinity it goes to 1/2?

OpenStudy (zarkon):

you want to show that \[(n^3+2n)/ 2n\le Mx^2\] for all x bigger than x_0 it is enough to show that I typed above

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

I'm still kind of confuse b/c I have the answer to a similar question but the answer is different

OpenStudy (anonymous):

it's for like (n2+1)/n

OpenStudy (zarkon):

\[(n^3+2n)/ 2n\le Mn^2\] n>n_0 (not x)

OpenStudy (anonymous):

and they choose values for m and n

OpenStudy (zarkon):

ok

OpenStudy (anonymous):

they reduce it n+(1/n) and then made it n+(1/n)<2n then they choose values for m=2 and k=1 and assume that n>k=1 so they have

OpenStudy (anonymous):

f(n)=(n^2+1)/n=n+(1/n)<2n = m*g(n)

OpenStudy (anonymous):

and that's the answer which does not make sense at all for me

OpenStudy (zarkon):

"f(n)=(n^2+1)/n=n+(1/n)<2n = m*g(n)" what don't you understand?

OpenStudy (anonymous):

I don't see how they got 2n because it was O(n)

OpenStudy (zarkon):

O(n) just means that \[(n^2+1)/n\le M n\] for some n (here n apparently greater than 1)

OpenStudy (zarkon):

if n>1 then 1/n<n so n+1/n<n+n=2n

OpenStudy (anonymous):

Ohh I see now so for mine n^2/2 + 1

OpenStudy (anonymous):

oh wait how do you know n is greater than 1

OpenStudy (zarkon):

usually you are looking at it as \[n\to \infty\] so assuming n>1 is no big deal

OpenStudy (anonymous):

okay so I simplified to 1/2 and that's <infinity

OpenStudy (anonymous):

I think taking the limit is a lot more simplify it's still the same interpretation i assume

OpenStudy (zarkon):

you could also do this... \[\frac{n^3+2n}{2n} =\frac{n^3}{2n}+\frac{2n}{2n}=\frac{n^2}{2}+1<n^2+n^2=2n^2=Mn^2\]

OpenStudy (zarkon):

both ways...using the limit or finding an upper bound works

OpenStudy (anonymous):

alright thanks so much for your help!!

OpenStudy (zarkon):

no problem